OpenStudy (anonymous):

given that \[\int\limits_{1}^{3} f(x) dx = -20 and \[\int\limits_{1}^{3} [f(x) +kx^2] dx= 6 find the value of k

OpenStudy (anonymous):

You can separate the second into two different integrals: \[6 =\int\limits\limits_{1}^{3}f(x)dx +\int\limits_{1}^{3}kx^2dx\] You already know the value of the first integral, so you can substitute the value you have in for that.

OpenStudy (anonymous):

\[\int\limits_{1}^{3} (-20+\frac{ k x^{3} }{ 3 }) dx = 6\]

OpenStudy (anonymous):

Not quite, you would substitute it like this: \[6=\int\limits_{1}^{3}f(x)dx + \int\limits_{1}^{3}kx^2dx\] \[\int\limits_{1}^{3}f(x)dx = -20\] \[6=(-20)+\int\limits_{1}^{3}kx^2dx\]

OpenStudy (anonymous):

ok, after that what? 20+ 6 and become 26 ?

OpenStudy (anonymous):

Correct. That then gives you: \[26=\int\limits_{1}^{3}kx^2dx\] Remember that constant values can be either kept inside or outside an integral that you are evaluating.

OpenStudy (anonymous):

[kx^3/3] is this correct so far?

OpenStudy (anonymous):

Yes, that is the indefinite integral of kx^2 (excluding the +c). What does that come out to be when we evaluate it as the definite integral from 1 to 3?

OpenStudy (anonymous):

wait, before that, I get this 27k/3 - k/3 = 26. 9k-k/3=26 9k-k=78 8k= 78 k-78/8 this is wrong right? idk how i can get this kind of ending. tell me my mistake

OpenStudy (anonymous):

Not quite. So where we are at is: \[26 = \int\limits_1^3kx^2dx = k\int\limits_1^3x^2dx\] \[26 = k * \frac{x^3}{3} |_1^3\] When we evaluate it at those x values, we get something like: \[26 = k *\left[ \frac{(3)^3}{3} \right]-\left[ \frac{(1)^3}{3} \right]\]

OpenStudy (anonymous):

wait, let me try first.

OpenStudy (anonymous):

wow, I get the actual answer, k=3 ! tq heril.:D